# Hydrostatic Newton-Cartan Membranes

### v1

@@ VI Conclusions
We have successfully used the formalism presented in Armas:2020 to construct the more general equilibrium partition function and, consequently, constitutive relations for NC membranes embedded in flat NC spacetime with both finite temperature and chemical potential. The equilibrium constraints imply that the temperature must be constant and the chemical potential must be related to the acceleration. The shape of the membrane in its equilibrium state will be determined by the generalized Young-Laplace equation (68). We wrote down this equation explicitely for a particular set of constitutive relations in the presence of a constant vorticity. We will leave the full study of solutions to this equation for future work. We will like to point out that in Anand2015 a relation between linear chemo-elasticity and linear poroelasticity was established, this means the presence of a mass chemical potential can be interpreted as having an underlying effective porous medium. This will allow us to eventually extend the range of applications of the resulting work in the context of finite temperature poroelastic membranes at hydrostatic equilibrium.
@@ Acknowledgement
This work was supported by Universidad Nacional Autónoma de México through UNAM-DGAPA’s posdoctoral fellowship program ”Programa de Becas Posdoctorales en la UNAM.”
Appendix A
In here we show the variation of each of the terms appearing on the equilibrium partition function
@@ A.3 Variations of order two scalars
$\delta S_{(1)}=2F_{(1)}\delta F_{(1)},$ (156)
@@
$+\int d^{d-1}\sigma\hat{e}\left[\left(4a_{1}^{(2)}F_{(1)}\hat{\upsilon}^{a}h^{bc}K_{ab}-\zeta^{(2)c}_{1}S_{(1)}\right)\delta\tau_{c}+\left(2a_{1}^{(2)}F_{(1)}h^{cd}\right)\delta K_{cd}\right].$
@@
$\delta S_{(2)}$ $=\left(-2h^{ac}h^{bd}h^{rs}K_{ar}K_{bs}\right)\delta\bar{h}_{cd}+\left(4\hat{\upsilon}^{a}h^{bc}h^{dr}K_{ad}K_{br}\right)\delta\tau_{c}+\left(2h^{ac}h^{bd}K_{ab}\right)\delta K_{cd},$ (158)
@@
$+\int d^{d-1}\sigma\hat{e}\left[\left(4a_{2}^{(2)}\hat{\upsilon}^{a}h^{bc}h^{dr}K_{ad}K_{br}-\zeta_{2}^{(2)c}S_{(2)}\right)\delta\tau_{c}+\left(2a_{2}^{(2)}h^{ac}h^{bd}K_{ab}\right)\delta K_{cd}\right].$
@@
$\delta S_{(3)}$ $=\left(-h^{cr}h^{ds}K_{ar}K_{bs}u^{a}u^{b}\right)\delta h_{cd}+\left(2\hat{\upsilon}^{a}h^{cr}K_{ab}K_{dr}u^{b}u^{d}-2u^{c}h^{dr}K_{ad}K_{br}u^{a}u^{b}\right)\delta\tau_{c}$ (160)
@@
$+\int d^{d-1}\sigma\hat{e}\left[\left(2a_{3}^{(2)}\hat{\upsilon}^{a}h^{cr}K_{ab}K_{dr}u^{b}u^{d}-2a_{3}^{(2)}u^{c}h^{dr}K_{ad}K_{br}u^{a}u^{b}-\zeta^{(2)c}_{3}S_{(3)}\right)\delta\tau_{c}\right]$
$+\int d^{d-1}\sigma\hat{e}\left[\left(2a_{3}^{(2)}u^{(c}h^{d)b}u^{a}K_{ab}\right)\delta K_{cd}\right].$
@@
$\delta\mathcal{F}_{S_{(4)}}$ $=\int d^{d-1}\sigma\hat{e}\left[\frac{1}{2}\left(\xi^{(2)cd}_{4}S_{(4)}\right)\delta\bar{h}_{cd}-\left(\zeta_{4}^{(2)c}S_{(4)}+4a_{4}^{(2)}F_{(2)}K_{ab}u^{a}u^{b}u^{c}\right)\delta\tau_{c}\right]$ (163)
@@
$\delta\mathcal{F}_{S_{(5)}}$ $=\frac{a^{(2)}_{5}}{2}\int d^{d-1}\sigma\hat{e}\left[\frac{\xi_{5}^{(2)cd}S_{(5)}}{a^{(2)}_{5}}+2a^{c}a^{d}+4a_{f}h^{cf}h^{dg}D_{d}\mu+2h^{fg}D_{f}a_{g}u^{c}u^{d}\right]\delta\bar{h}_{cd}$ (165)
$-\int d^{d-1}\sigma\hat{e}\left[\left(\zeta_{5}^{(2)c}S_{(5)}+4a^{(2)}_{5}\hat{\upsilon}^{(f}h^{g)c}a_{f}D_{g}\mu+2a^{(2)}_{5}u^{c}\left(\mu+\frac{1}{2}\bar{u}^{2}\right)h^{fg}D_{f}D_{g}\mu\right)\delta\tau_{c}\right].$
@@
$\delta\mathcal{F}_{S_{(6)}}$ $=\int\frac{d^{d-1}\sigma\hat{e}}{2}\left[\xi^{(2)cd}_{6}S_{(6)}+a^{(2)}_{6}\left(2K_{ab}u^{a}h^{bc}h^{fd}D_{f}\mu+h^{bf}D_{f}\left(K_{ab}u^{a}\right)u^{c}u^{d}\right)\right]\delta\bar{h}_{cd}$ (167)
$-\int d^{d-1}\sigma\hat{e}\left(\zeta^{(2)c}_{6}S_{(6)}+a^{(2)}_{(6)}K_{fg}a^{f}u^{g}u^{c}+2a^{(2)}_{6}K_{ab}u^{a}\hat{\upsilon}^{(b}h^{f)c}D_{f}\mu\right)\delta\tau_{c}$
$+\int d^{d-1}\sigma\hat{e}\left[-h^{bf}D_{f}\left(a^{(2)}_{6}K_{ab}u^{a}\right)\left(\mu+\frac{1}{2}\bar{u}^{2}\right)u^{a}\delta\tau_{a}+\left(a^{(2)}_{6}a^{c}u^{d}\right)\delta K_{cd}\right].$
@@
$=\left[\bar{u}_{a}\bar{h}_{bc}\Omega^{ab}-D_{c}\left(\mu+\frac{1}{2}\bar{u}^{2}\right)\right]\Omega^{cd}\delta\tau_{d}-\left(\Omega^{af}\Omega^{bg}\bar{h}_{ab}\bar{h}_{fg}\right)u^{c}\delta\tau_{c}$ (169)
@@
$-\int d^{d-1}\sigma\hat{e}\left[\zeta^{(2)d}_{7}S_{(7)}-a^{(2)}_{(7)}\bar{u}_{a}\bar{h}_{bc}\Omega^{ab}\Omega^{cd}+a^{(2)}_{7}D_{c}\left(\mu+\frac{1}{2}\bar{u}^{2}\right)\Omega^{cd}\right]\delta\tau_{d}$
$+\int d^{d-1}\sigma\hat{e}\left[D_{f}\left(a^{(2)}_{(7)}\bar{h}_{bc}\Omega^{cf}u^{b}\right)-a^{(2)}_{(7)}\left(\Omega^{af}\Omega^{bg}\bar{h}_{ab}\bar{h}_{fg}\right)\right]u^{c}\delta\tau_{c}.$

### v2

@@ 6 Conclusions
We have successfully used the formalism presented in [3] to construct the equilibrium partition function and the constitutive relations for co-dimension one NC membranes embedded in flat NC spacetime with both finite temperature and chemical potential up to second order in a derivative expansion. The equilibrium constraints imply that the temperature must be constant and the chemical potential must be related to the acceleration. The shape of the membrane in its equilibrium state will be determined by the generalized Young-Laplace equation (69). We wrote down this equation explicitly for a particular set of constitutive relations in the presence of constant vorticity. We would like to point out that in [1] a relation between linear chemo-elasticity and linear poroelasticity was established, which means the presence of a mass chemical potential can be interpreted as having an underlying effective porous medium. This will allow us to eventually extend the range of applications of the resulting work in the context of finite temperature poroelastic membranes at hydrostatic equilibrium.
As happens in the case of a static closed membrane [5], spheres, tori, and biconcave discoids are solutions to the Young-Laplace equations when $\omega\neq 0$, as long as the elastic moduli and spontaneous curvature are linear function of the chemical potential $\mu$ and specific relations are met by the surface tension and pressure with the elastic moduli and spontaneous curvature parameters. Additionally, it is possible to find numerically a family of axisymmetric solutions with material conditions that allow us to continuously connect them to the $\omega=0$ biconcave disc analytic solution.
@@ Acknowledgement
This work was supported by Universidad Nacional Autónoma de México through UNAM-DGAPA’s posdoctoral fellowship program "Programa de Becas Posdoctorales en la UNAM."
Appendix A Appendix
In here we show the variation of each of the terms appearing in the equilibrium partition function
@@ A.1 Variation of $\chi$
The contribution of this variation to the currents is given by
$\left(\mathcal{T}_{S}\right)^{c}_{\chi}$ $=-\chi\hat{\upsilon}^{c}-\left(\epsilon+\chi+\frac{n}{s}\bar{u}^{2}\right)u^{c},$ (185)
$\left(\mathcal{T}_{S}\right)^{ab}_{\chi}$ $=\chi h^{ab}+nu^{a}u^{b},$ (186)
$D^{cd}_{\chi}$ $=0.$ (187)
@@ A.2 Variations of order one scalars
The contribution of this variation to the currents is given by
$\left(\mathcal{T}_{S}\right)^{c}_{F_{(1)}}$ $=-a_{1}^{(1)}K\hat{\upsilon}^{c}-\left[s_{1}^{(1)}T+n_{1}^{(1)}\left(\mu+\frac{1}{2}\bar{u}^{2}\right)\right]Ku^{c}+2a^{(1)}_{1}\hat{\upsilon}^{a}h^{bc}K_{ab},$ (190)
$\left(\mathcal{T}_{S}\right)^{cd}_{F_{(1)}}$ $=a^{(1)}_{1}Kh^{cd}+n^{(1)}_{1}Ku^{c}u^{d}-2a^{(1)}_{1}h^{ac}h^{bd}K_{ab},$ (191)
$D^{cd}_{F_{(1)}}$ $=a_{1}^{(1)}h^{cd}.$ (192)
@@
The contribution of this variation to the currents is given by
$\left(\mathcal{T}_{S}\right)^{c}_{F_{(2)}}$ $=-a_{2}^{(1)}\left(a\cdot n\right)\hat{\upsilon}^{c}-\left[s_{2}^{(1)}T+n_{2}^{(1)}\left(\mu+\frac{1}{2}\bar{u}^{2}\right)+2a^{(1)}_{2}\right]\left(a\cdot n\right)u^{c},$ (195)
$\left(\mathcal{T}_{S}\right)^{cd}_{F_{(2)}}$ $=0,$ (196)
$\mathcal{D}^{cd}_{F_{(2)}}$ $=a^{(1)}_{2}u^{c}u^{d}\,.$ (197)
@@ A.3 Variations of order two scalars
$\delta S_{(1)}=2F_{(1)}\delta F_{(1)}.$ (198)
Hence, it will contribute to the variation of the equilibrium partition function as
@@
$+\int d^{d-1}\sigma\hat{e}\left[\left(4a_{1}^{(2)}F_{(1)}\hat{\upsilon}^{a}h^{bc}K_{ab}-\zeta^{(2)c}_{1}S_{(1)}\right)\delta\tau_{c}\right$.
$\left.+\left(2a_{1}^{(2)}F_{(1)}h^{cd}\right)\delta K_{cd}\right].$
The contribution of this variation to the currents is given by
$\left(\mathcal{T}_{S}\right)^{c}_{S_{(1)}}$ $=-a^{(2)}_{1}K^{2}\hat{\upsilon}^{c}-\left[s^{(2)}_{1}T+n^{(2)}_{1}\left(\mu+\frac{1}{2}\bar{u}^{2}\right)\right]K^{2}u^{c}+4a^{(2)}_{1}K\hat{\upsilon}^{a}h^{bc}K_{ab},$ (200)
$\left(\mathcal{T}_{S}\right)^{cd}_{S_{(1)}}$ $=a^{(1)}_{2}K^{2}h^{cd}+n^{(2)}_{1}K^{2}u^{c}u^{d}-4a^{(2)}_{1}Kh^{ac}h^{bd}K_{ab},$ (201)
$\mathcal{D}^{cd}_{S_{(1)}}$ $=2a^{(2)}_{1}Kh^{cd}.$ (202)
@@
$\delta S_{(2)}$ $=\left(-2h^{ac}h^{bd}h^{rs}K_{ar}K_{bs}\right)\delta\bar{h}_{cd}+\left(4\hat{\upsilon}^{a}h^{bc}h^{dr}K_{ad}K_{br}\right)\delta\tau_{c}$
$+\left(2h^{ac}h^{bd}K_{ab}\right)\delta K_{cd},$ (203)
@@
$+\int d^{d-1}\sigma\hat{e}\left[\left(4a_{2}^{(2)}\hat{\upsilon}^{a}h^{bc}h^{dr}K_{ad}K_{br}-\zeta_{2}^{(2)c}S_{(2)}\right)\delta\tau_{c}\right$.
$\left.+\left(2a_{2}^{(2)}h^{ac}h^{bd}K_{ab}\right)\delta K_{cd}\right].$ (205)
The contribution of this variation to the currents is given by
$\left(\mathcal{T}_{S}\right)^{c}_{S_{(2)}}$ $=-a^{(2)}_{2}\left(K\cdot K\right)\hat{\upsilon}^{c}-\left[s^{(2)}_{2}T+n^{(2)}_{2}\left(\mu+\frac{1}{2}\bar{u}^{2}\right)\right]\left(K\cdot K\right)u^{c}$ (206)
$+4a^{(2)}_{2}\hat{\upsilon}^{a}K_{ad}h^{dr}K_{rb}h^{bc},$
$\left(\mathcal{T}_{S}\right)^{cd}_{S_{(2)}}$ $={a^{(2)}_{2}\left(K\cdot K\right)h^{cd}+n^{(2)}_{2}\left(K\cdot K\right)u^{c}u^{d}-4a^{(2)}_{2}K_{ra}h^{rs}K_{sb}h^{ac}h^{bd},}$ (207)
$\mathcal{D}^{cd}_{S_{(2)}}$ $=2a^{(2)}_{2}h^{ac}h^{bd}K_{ab}.$ (208)
@@
$\delta S_{(3)}$ $=\left(-h^{cr}h^{ds}K_{ar}K_{bs}u^{a}u^{b}\right)\delta h_{cd}$ (209)
$+\left(2\hat{\upsilon}^{a}h^{cr}K_{ab}K_{dr}u^{b}u^{d}-2u^{c}h^{dr}K_{ad}K_{br}u^{a}u^{b}\right)\delta\tau_{c}$
@@
[…]
$\left.\left.-\zeta^{(2)c}_{3}S_{(3)}\right)\delta\tau_{c}\right]+\int d^{d-1}\sigma\hat{e}\left[\left(2a_{3}^{(2)}u^{(c}h^{d)b}u^{a}K_{ab}\right)\delta K_{cd}\right].$
The contribution of this variation to the currents is given by
$\left(\mathcal{T}_{S}\right)^{c}_{S_{(3)}}$ $=-a^{(2)}_{3}\left(n\cdot\Omega\right)^{2}\hat{\upsilon}^{c}-\left[s_{3}^{(2)}T+n^{(2)}_{3}\left(\mu+\frac{1}{2}\bar{u}^{2}\right)\right]\left(n\cdot\Omega\right)^{2}u^{c}$ (211)
$+2a^{(2)}_{3}\hat{\upsilon}^{a}h^{cr}K_{ab}K_{dr}u^{b}u^{d}-2a^{(2)}_{3}u^{c}h^{dr}K_{ad}K_{br}u^{a}u^{b},$
$\left(\mathcal{T}_{S}\right)^{cd}_{S_{(3)}}$ $=a^{(2)}_{3}\left(n\cdot\Omega\right)^{2}h^{cd}+n_{3}^{(2)}\left(n\cdot\Omega\right)^{2}u^{c}u^{d}-2a_{3}^{(2)}h^{cr}h^{ds}K_{ar}K_{bs}u^{a}u^{b},$ (212)
$\mathcal{D}^{cd}_{S_{(3)}}$ $=2a^{(2)}_{3}u^{(c}h^{d)b}u^{a}K_{ab}.$ (213)
@@
$\delta\mathcal{F}_{S_{(4)}}$ $=\int d^{d-1}\sigma\hat{e}\left[\frac{1}{2}\left(\xi^{(2)cd}_{4}S_{(4)}\right)\delta\bar{h}_{cd}\right$. (215)
$\left.-\left(\zeta_{4}^{(2)c}S_{(4)}+4a_{4}^{(2)}F_{(2)}K_{ab}u^{a}u^{b}u^{c}\right)\delta\tau_{c}\right]$
@@
The contribution of this variation to the currents is given by
$\left(\mathcal{T}_{S}\right)^{c}_{S_{(4)}}$ $=-a^{(2)}_{4}\left(a\cdot n\right)^{2}\hat{\upsilon}^{c}-\left[s_{4}^{(2)}T+n^{(2)}_{4}\left(\mu+\frac{1}{2}\bar{u}^{2}\right)\right]\left(a\cdot n\right)^{2}u^{c}$ (216)
$-4a^{(2)}_{4}\left(a\cdot n\right)^{2}u^{c},$
$\left(\mathcal{T}_{S}\right)^{cd}_{S_{(4)}}$ $=a^{(2)}_{4}\left(a\cdot n\right)^{2}h^{cd}+n^{(2)}_{4}\left(a\cdot n\right)^{2}u^{c}u^{d},$ (217)
$\mathcal{D}^{cd}_{S_{(4)}}$ $=2\left(a\cdot n\right)a^{(2)}_{4}u^{c}u^{d}.$ (218)
@@
$\delta\mathcal{F}_{S_{(5)}}$ $=\frac{a^{(2)}_{5}}{2}\int d^{d-1}\sigma\hat{e}\left[\frac{\xi_{5}^{(2)cd}S_{(5)}}{a^{(2)}_{5}}+2a^{c}a^{d}+4a_{f}h^{cf}h^{dg}D_{d}\mu\right$. (220)
$\left.+2h^{fg}D_{f}\bar{a}_{g}u^{c}u^{d}\right]\delta\bar{h}_{cd}$
[…]
$\left.\left.+2a^{(2)}_{5}u^{c}\left(\mu+\frac{1}{2}\bar{u}^{2}\right)h^{fg}D_{f}D_{g}\mu\right)\delta\tau_{c}\right].$
The contribution of this variation to the currents is given by
$\left(\mathcal{T}_{S}\right)^{c}_{S_{(5)}}$ $=-a^{(2)}_{5}\bar{a}^{2}\hat{\upsilon}^{c}-\left[s_{5}^{(2)}T+n^{(2)}_{5}\left(\mu+\frac{1}{2}\bar{u}^{2}\right)\right]\bar{a}^{2}u^{c}$ (221)
$-4a^{(2)}_{5}\hat{\upsilon}^{(f}h^{g)c}a_{f}D_{g}\mu-2a^{(2)}_{5}u^{c}\left(\mu+\frac{1}{2}\bar{u}^{2}\right)h^{fg}D_{f}D_{g}\mu,$
$\left(\mathcal{T}_{S}\right)^{cd}_{S_{(5)}}$ $=a^{(2)}_{5}\bar{a}^{2}h^{cd}+n^{(2)}_{5}\bar{a}^{2}u^{c}u^{d}+2a^{(2)}_{5}a^{c}a^{d}+4a^{(2)}_{5}\bar{a}_{f}h^{f(c}h^{d)g}D_{g}\mu$ (222)
$+2a^{(2)}_{5}h^{fg}\left(D_{f}\bar{a}_{g}\right)u^{c}u^{d},$
$\mathcal{D}^{cd}_{S_{(5)}}$ $=0.$ (223)
@@
[…]
$\left.\left.+h^{bf}D_{f}\left(K_{ab}u^{a}\right)u^{c}u^{d}\right)\right]\delta\bar{h}_{cd}$
$-\int d^{d-1}\sigma\hat{e}\left(\zeta^{(2)c}_{6}S_{(6)}+a^{(2)}_{6}K_{fg}a^{f}u^{g}u^{c}+2a^{(2)}_{6}K_{ab}u^{a}\hat{\upsilon}^{(b}h^{f)c}D_{f}\mu\right)\delta\tau_{c}$
$+\int d^{d-1}\sigma\hat{e}\left[-h^{bf}D_{f}\left(a^{(2)}_{6}K_{ab}u^{a}\right)\left(\mu+\frac{1}{2}\bar{u}^{2}\right)u^{c}\delta\tau_{c}\right$.
$\left.+\left(a^{(2)}_{6}a^{c}u^{d}\right)\delta K_{cd}\right].$
The contribution of this variation to the currents is given by
$\left(\mathcal{T}_{S}\right)^{c}_{S_{(6)}}$ $=-a^{(2)}_{6}\left(\bar{a}\cdot\left(n\cdot\Omega\right)\right)\hat{\upsilon}^{c}-\left[s_{6}^{(2)}T+n^{(2)}_{6}\left(\mu+\frac{1}{2}\bar{u}^{2}\right)\right]\left(\bar{a}\cdot\left(n\cdot\Omega\right)\right)u^{c}$
$-a^{(2)}_{6}K_{fg}a^{f}u^{g}u^{c}-2a^{(2)}_{6}K_{ab}u^{a}\hat{\upsilon}^{(b}h^{f)c}D_{f}\mu$
$-h^{bf}D_{f}\left(a^{(2)}_{6}K_{ab}u^{b}\right)\left(\mu+\frac{1}{2}\bar{u}^{2}\right)u^{c},$ (226)
$\left(\mathcal{T}_{S}\right)^{cd}_{S_{(6)}}$ $=a^{(2)}_{6}\left(\bar{a}\cdot\left(n\cdot\Omega\right)\right)h^{cd}+n^{(2)}_{6}\left(\bar{a}\cdot\left(n\cdot\Omega\right)\right)u^{c}u^{d}$
$+2a^{(2)}_{6}K_{ab}u^{a}h^{bc}h^{fd}D_{f}\mu+a^{(2)}_{6}h^{bf}D_{f}\left(K_{ab}u^{a}\right)u^{c}u^{d},$ (227)
$\mathcal{D}^{cd}_{S_{(6)}}$ $=a^{(2)}_{6}a^{(c}u^{d)}.$ (228)
@@
$=\left[\bar{u}_{a}\bar{h}_{bc}\Omega^{ab}-D_{c}\left(\mu+\frac{1}{2}\bar{u}^{2}\right)\right]\Omega^{cd}\delta\tau_{d}-\left(\Omega^{af}\Omega^{bg}\bar{h}_{ab}\bar{h}_{fg}\right)u^{c}\delta\tau_{c}$
@@
$-\int d^{d-1}\sigma\hat{e}\left[\zeta^{(2)d}_{7}S_{(7)}-a^{(2)}_{7}\bar{u}_{a}\bar{h}_{bc}\Omega^{ab}\Omega^{cd}\right$.
$\left.+a^{(2)}_{7}D_{c}\left(\mu+\frac{1}{2}\bar{u}^{2}\right)\Omega^{cd}\right]\delta\tau_{d}$
$+\int d^{d-1}\sigma\hat{e}\left[D_{f}\left(a^{(2)}_{7}\bar{h}_{bd}\Omega^{cd}u^{b}\right)-a^{(2)}_{7}\left(\Omega^{af}\Omega^{bg}\bar{h}_{ab}\bar{h}_{fg}\right)\right]u^{c}\delta\tau_{c}.$
The contribution of this variation to the currents is given by
$\left(\mathcal{T}_{S}\right)^{c}_{S_{(7)}}$ $=-a^{(2)}_{7}\bar{\Omega}^{2}\hat{\upsilon}^{c}-\left[s_{6}^{(7)}T+n^{(2)}_{7}\left(\mu+\frac{1}{2}\bar{u}^{2}\right)\right]\bar{\Omega}^{2}u^{c}$ (231)
$+a^{(2)}_{7}\bar{u}_{a}\bar{h}_{bd}\Omega^{ab}\Omega^{dc}-a^{(2)}_{7}D_{d}\left(\mu+\frac{1}{2}\bar{u}^{2}\right)\Omega^{dc}$
$+D_{f}\left(a^{(2)}_{7}\bar{h}_{bd}\Omega^{df}u^{b}\right)u^{c}-a^{(2)}_{7}\left(\Omega^{af}\Omega^{bg}\bar{h}_{ab}\bar{h}_{fg}\right)u^{c},$
$\left(\mathcal{T}_{S}\right)^{cd}_{S_{(7)}}$ $=a^{(2)}_{7}\bar{\Omega}^{2}h^{cd}+n^{(2)}_{7}\bar{\Omega}^{2}u^{c}u^{d}-2a^{(2)}_{7}\Omega^{f(c}\Omega^{d)g}\bar{h}_{fg}-2D_{b}\left(a^{(2)}_{7}\Omega^{b(c}u^{d)}\right),$ (232)
$\mathcal{D}^{cd}_{S_{(7)}}$ $=0.$ (233)
Appendix B Appendix
Figure 2 shows the plot of the aspect ratio $\mathcal{A}$ of numerical solutions to (5), vs the dimensionless vorticity $\hat{\omega}_{\bar{\kappa}}^{2}$, which can take negative values. Similar to figure 1, horizontal dashed lines determine intervals of $\mathcal{A}$ that correspond to different characteristic shapes, the same ones shown in figure 1.
Figure 2: Plot of $\mathcal{A}$ vs $\hat{\omega}_{\bar{\kappa}}^{2}$ for a series of values of $rK_{0}^{(0)}$.